# Topological phase transition through tunable nearest-neighbor interactions in a one-dimensional lattice

### v1

@@ IV Conclusion
We have numerically investigated the ground state properties of a one-dimensional lattice of hardcore bosons or spinless fermions with tunable NN interactions. While the non-interacting counterpart of the system is a trivial gapless phase, we have showed that by allowing attractive and repulsive interactions on the alternate NN bonds of the lattice, a topological phase can be established. We have also showed that by tuning the NN interaction appropriately, a topological to trivial phase transition occurs at a critical point where the interactions vanish. Apart from this, we have also obtained a symmetry broken phase for stronger interaction strengths. While we have revealed that the bulk of the topological and the trivial phases exhibit finite bond ordering, the symmetry broken phase exhibits a CDW order. We have obtained the phase diagram depicting these phases and the transitions between them at half-filling with tunable NN interactions. Furthermore, by extending our analysis beyond half-filling, we have explored the phase diagram at all densities and identified both single-particle and bound-pair excitations at incommensurate fillings, leading to the superfluid (SF) and pair-superfluid (PSF) phases. We have also provided analytical arguments for specific scenarios in support of the numerical results. Finally, we discussed the experimental feasibility of realizing the model under consideration in quantum simulation platforms, such as Rydberg atom and ultracold atomic set ups.
V Acknowledgements
@@
Appendix A Degenerate perturbation theory for the anisotropic limits
@@
In the space of degenerate ground states described above, we see that the $V_{1}$ term in Eq. (3) effectively acts as $-(V_{1}/4)\tau_{j}^{z}\tau_{j+1}^{z}$. For instance, acting on the state $|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle=|\uparrow_{i}\downarrow_{i+1}\uparrow_{i+2}\downarrow_{i+3}\rangle$, the term $V_{1}S_{i+1}^{z}S_{i+2}^{z}$ gives $(-V_{1}/4)|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle$.
@@
$-~{}2t~{}(S_{i}^{x}S_{i+1}^{x}~{}+~{}S_{i}^{y}S_{i+1}^{y})~{}=~{}-~{}t~{}(\sigma_{i}^{+}\sigma_{i+1}^{-}~{}+~{}\sigma_{i}^{-}\sigma_{i+1}^{+})$ (28)
flips the state $\tau_{j}^{z}=1$ to $\tau_{j}^{z}=-1$, and vice versa. Hence this term effectively acts as $-t~{}\tau_{j}^{x}$.
@@
$-~{}2t~{}(S_{i-1}^{x}S_{i}^{x}~{}+~{}S_{i-1}^{y}S_{i}^{y})~{}=~{}-~{}t~{}(\sigma_{i-1}^{+}\sigma_{i}^{-}~{}+~{}\sigma_{i-1}^{-}\sigma_{i}^{+}),$ (29)
where $i$ is even, does not have any matrix elements between states in the space of degenerate ground states. But this term contributes to second order as follows. Acting on the state $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=1\rangle=|\uparrow_{i-2}\downarrow_{i-1}\uparrow_{i}\downarrow_{i+1}\rangle$, the term in Eq. (29) gives $-t|\uparrow_{i-2}\uparrow_{i-1}\downarrow_{i}\downarrow_{i+1}\rangle$ which is an excited state whose energy is $V_{2}$ larger than the energy of the degenerate ground states. Another application of Eq. (29) brings the excited state back to $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=1\rangle$, with a coefficient equal to $(-t)^{2}/(-V_{2})=-t^{2}/V_{2}$ as given by second-order perturbation theory; the energy denominator equal to $-V_{2}$ comes from $E_{ground}-E_{excited}$. A similar terms appears for the state $|\tau_{j-1}^{z}=-1,\tau_{j}^{z}=-1\rangle$. However, no such terms appear for the states $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=-1\rangle$ and $|\tau_{j-1}^{z}=-1,\tau_{j}^{z}=1\rangle$ since Eq. (29) gives zero on these states. Hence, Eq. (29) gives rise to a term in the effective Hamiltonian given by $(-t^{2}/(2V_{2}))(\mathbb{I}~{}+~{}\tau_{j-1}^{z}\tau_{j}^{z})$.
@@
This describes the transverse field Ising model. It is well-known that this model has a second-order phase transition when the coefficients of the two terms in Eq. (30) are equal in magnitude, i.e., when $t=|(V_{1}/4)+(t^{2}/(2V_{2}))|$. For $V_{1}<0$, this gives the line
@@ Appendix B Boundaries close to $\rho=1$ and $\rho=0$
In this Appendix, we will consider the Hamiltonian in Eq. (1) with $V_{1}=-V_{2}=V$, and add a chemical potential term $-\mu\sum_{i}n_{i}$. We will derive the forms of the boundaries close to the completely filled and completely empty states that we see in Fig. 12 for open boundary conditions and Fig. 16 for periodic boundary conditions.
To understand Fig. 16, we will consider an $L$-site system with periodic boundary conditions so that momentum eigenstates can be defined. The fully occupied state in which the number of particles is $L$ is an eigenstate of Eq. (1) with energy equal to $E_{0}=-\mu L$. Next, we consider a one-hole state in which the number of particles is $L-1$. Defining a state with a hole located at site $j$ as $|j\rangle$, we find that
@@
We will now study the form of a two-hole bound state when $|V|$ may be of the same order as $t$. A two-hole state will have sites $n_{1}$ and $n_{2}$ empty, where $n_{1}<n_{2}$; we will denote this state as $|n_{1},n_{2}\rangle$. (In this discussion, it will be convenient to consider an infinitely large system so that we do not have to worry about boundary conditions). We will now consider a bound state which has a center-of-mass momentum $k$; this lies in the range $[-2\pi,2\pi]$. The wave function of such a state can be written as
@@
Hence, if there is a solution of Eqs. (40) with energy $E$, there must be another solution with energy $-E$.
For every value of $k$, we can numerically solve Eqs. (40) and see if there are bound state solutions; these must satisfy $|\alpha(m)|$ and $|\beta(m)|\to 0$ as $m\to\infty$. If a bound state exists, we denote its energy as $E_{2}(k)$. (Then there will be another bound state with energy $-E_{2}(k)$ as discussed above). As $|V|$ is increased from zero, we find that bound states first appears at $k=0$. Further, if bound states are present for a range of values of $k$, its energy is the lowest at $k=0$. We will therefore concentrate on the case $k=0$ henceforth. Eqs. (42) then take the form
@@
$-2t~{}[\alpha(m+1)~{}+~{}\alpha(m-1)]~{}=~{}E_{2}(0)~{}\alpha(m)~{}~{}{\rm for}~{}~{}m\geq 3.$ (45)
@@
As argued above, there must be another bound state at $k=0$ with the energy $-E_{2}(k=0)$.
Since a bound state must have $\lambda>0$, Eq. (48) implies that a bound state appears only if $|V|>2\sqrt{2}t$. Since the right hand side of Eq. (49) is always less than $-2t$, we see that the lowest energy excitations around the fully occupied state are two-hole bound states rather than one-hole states, if $|V|>2\sqrt{2}t$.
@@
where $|n\rangle$ denotes a state in which only site $n$ is empty, and $n=1$ denotes the leftmost site of the system. From Eq. (1), we find that the diagonal term of the Hamiltonian has a value which is $V/2$ less at $n=1$ as compared to all other values of $n\geq 2$. Next, including the hopping term proportional to $-t$, we effectively obtain a tight-binding model with a negative on-site potential $-V/2$ at site 1 and hopping amplitude $-t$ between neighboring sites $n$ and $n+1$. Eq. (52) and the eigenvalue condition $H|\psi\rangle_{B}=E_{B}|\psi\rangle_{B}$ lead to the equations

### v2

@@ V Conclusion
We have numerically investigated the ground state properties of a one-dimensional lattice of hardcore bosons or spinless fermions with tunable NN interactions. While the non-interacting counterpart of the system is a trivial gapless phase, we have shown that by allowing attractive and repulsive interactions on the alternate NN bonds of the lattice, a topological phase can be established. We have also shown that by tuning the NN interaction appropriately, a topological to trivial phase transition occurs at a critical point where the interactions vanish. Apart from this, we have also obtained a symmetry broken phase for stronger interaction strengths. While we have revealed that the bulk of the topological and the trivial phases exhibit finite bond ordering, the symmetry broken phase exhibits a CDW order. We have obtained the phase diagram depicting these phases and the transitions between them at half-filling with tunable NN interactions. Furthermore, by extending our analysis beyond half-filling, we have explored the phase diagram at all densities and identified both single-particle and bound-pair excitations at incommensurate fillings, leading to the superfluid (SF) and pair-superfluid (PSF) phases. We have also provided analytical arguments for specific scenarios in support of the numerical results. Finally, we discussed the experimental feasibility of realizing the model under consideration in quantum simulation platforms, such as Rydberg atom and ultracold atomic set ups. To complete the study, we discussed the experimental feasibility of realizing the model under consideration in quantum simulation platforms, such as Rydberg atom and ultracold atomic setups. Finally, by systematically exploring all possible combinations of attractive and repulsive NN interactions, we presented the full interaction-driven phase diagram, offering a comprehensive understanding of the emergent quantum phases and their transitions.
VI Acknowledgements
@@
Appendix A Perturbation theory in the limit $V_{2}\to\infty$
@@
In the space of degenerate ground states described above, we see that the $V_{1}$ term in Eq. (3) effectively acts as $(V_{1}/4)\tau_{j}^{z}\tau_{j+1}^{z}$. For instance, acting on the state $|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle=|\uparrow_{i}\downarrow_{i+1}\uparrow_{i+2}\downarrow_{i+3}\rangle$, the term $V_{1}S_{i+1}^{z}S_{i+2}^{z}$ gives $(-V_{1}/4)|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle$.
@@
$2t~(S_{i}^{x}S_{i+1}^{x}~+~S_{i}^{y}S_{i+1}^{y})~=~-~t~(\sigma_{i}^{+}\sigma_{i+1}^{-}~+~\sigma_{i}^{-}\sigma_{i+1}^{+})$ (29)
flips the state $\tau_{j}^{z}=1$ to $\tau_{j}^{z}=-1$, and vice versa. Hence this term effectively acts as $t~\tau_{j}^{x}$.
@@
$2t~(S_{i-1}^{x}S_{i}^{x}~+~S_{i-1}^{y}S_{i}^{y})~=~-~t~(\sigma_{i-1}^{+}\sigma_{i}^{-}~+~\sigma_{i-1}^{-}\sigma_{i}^{+}),$ (30)
where $i$ is even, does not have any matrix elements between states in the space of degenerate ground states. But this term contributes to second order as follows. Acting on the state $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=1\rangle=|\uparrow_{i-2}\downarrow_{i-1}\uparrow_{i}\downarrow_{i+1}\rangle$, the term in Eq. (30) gives $t|\uparrow_{i-2}\uparrow_{i-1}\downarrow_{i}\downarrow_{i+1}\rangle$ which is an excited state whose energy is $V_{2}$ larger than the energy of the degenerate ground states. Another application of Eq. (30) brings the excited state back to $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=1\rangle$, with a coefficient equal to $(-t)^{2}/(-V_{2})=-t^{2}/V_{2}$ as given by second-order perturbation theory; the energy denominator equal to $V_{2}$ comes from $E_{ground}-E_{excited}$. A similar terms appears for the state $|\tau_{j-1}^{z}=-1,\tau_{j}^{z}=-1\rangle$. However, no such terms appear for the states $|\tau_{j-1}^{z}=1,\tau_{j}^{z}=-1\rangle$ and $|\tau_{j-1}^{z}=-1,\tau_{j}^{z}=1\rangle$ since Eq. (30) gives zero on these states. Hence, Eq. (30) gives rise to a term in the effective Hamiltonian given by $(-t^{2}/(2V_{2}))(\mathbb{I}~+~\tau_{j-1}^{z}\tau_{j}^{z})$.
@@
This describes the transverse field Ising model (called the quantum Ising model). It is well-known that this model has a second-order phase transition when the coefficients of the two terms in Eq. (31) are equal in magnitude, i.e., when $t=|(V_{1}/4)+(t^{2}/(2V_{2}))|$. For $V_{1}<0$, this gives the line
@@ Appendix B Boundaries close to $\rho=1$ and $\rho=0$
In this Appendix, we will consider the Hamiltonian in Eq. (1) with $V_{1}=-V_{2}=V$, and add a chemical potential term $\mu\sum_{i}n_{i}$. We will derive the forms of the boundaries close to the completely filled and completely empty states that we see in Fig. 14 for open boundary conditions and Fig. 18 for periodic boundary conditions.
To understand Fig. 18, we will consider an $L$-site system with periodic boundary conditions so that momentum eigenstates can be defined. The fully occupied state in which the number of particles is $L$ is an eigenstate of Eq. (1) with energy equal to $E_{0}=-\mu L$. Next, we consider a one-hole state in which the number of particles is L1L-1. Defining a state with a hole located at site $j$ as $|j\rangle$, we find that
@@
We will now study the form of a two-hole bound state when $|V|$ may be of the same order as $t$. A two-hole state will have sites $n_{1}$ and $n_{2}$ empty, where $n_{1}<n_{2}$; we will denote this state as $|n_{1},n_{2}\rangle$. (In this discussion, it will be convenient to consider an infinitely large system so that we do not have to worry about boundary conditions). We will now consider a bound state which has a center-of-mass momentum $k$; this lies in the range $2\pi,2\pi]$. The wave function of such a state can be written as
@@
Hence, if there is a solution of Eqs. (41) with energy $E$, there must be another solution with energy E-E.
For every value of $k$, we can numerically solve Eqs. (41) and see if there are bound state solutions; these must satisfy $|\alpha(m)|$ and $|\beta(m)|\to 0$ as $m\to\infty$. If a bound state exists, we denote its energy as $E_{2}(k)$. (Then there will be another bound state with energy $E_{2}(k)$ as discussed above). As $|V|$ is increased from zero, we find that bound states first appears at $k=0$. Further, if bound states are present for a range of values of $k$, its energy is the lowest at $k=0$. We will therefore concentrate on the case $k=0$ henceforth. Eqs. (43) then take the form
@@
$2t~[\alpha(m+1)~+~\alpha(m-1)]~=~E_{2}(0)~\alpha(m)~~{\rm for}~~m\geq 3.$ (46)
@@
As argued above, there must be another bound state at $k=0$ with the energy $E_{2}(k=0)$.
Since a bound state must have $\lambda>0$, Eq. (49) implies that a bound state appears only if $|V|>2\sqrt{2}t$. Since the right hand side of Eq. (50) is always less than 2t-2t, we see that the lowest energy excitations around the fully occupied state are two-hole bound states rather than one-hole states, if $|V|>2\sqrt{2}t$.
@@
where $|n\rangle$ denotes a state in which only site $n$ is empty, and $n=1$ denotes the leftmost site of the system. From Eq. (1), we find that the diagonal term of the Hamiltonian has a value which is $V/2$ less at $n=1$ as compared to all other values of $n\geq 2$. Next, including the hopping term proportional to t-t, we effectively obtain a tight-binding model with a negative on-site potential V/2-V/2 at site 1 and hopping amplitude t-t between neighboring sites $n$ and $n+1$. Eq. (53) and the eigenvalue condition $H|\psi\rangle_{B}=E_{B}|\psi\rangle_{B}$ lead to the equations
@@
Appendix C Perturbation theory in the limit $V_{1}\to$-$\infty$
In this Appendix we will provide a derivation of the transition line separating the phases CDW-II and PS in Fig. 21. The analysis will proceed in the same way as in Appendix A. Given the Hamiltonian in Eq. (5), we will assume that $V_{1}<0$, with $V_{1}\gg t$ and $|V_{2}|$, and we will go up to second order in perturbation theory. We begin by considering only the terms with $V_{1}$ which constitutes the unperturbed Hamiltonian. The ground state of this part of the Hamiltonian in Eq. (3) has a large degeneracy: for every odd value of $i$, the state at sites $i$ and $i+1$ must be either $|1_{i}1_{i+1}\rangle$ or $|0_{i}0_{i+1}\rangle$, where $1_{i}$ and $0_{i}$ denote the possible fermion occupation numbers $n_{i}$ at site $i$. We then define a new lattice with sites labeled as j=(i1)/2j=(i-1)/2 and spin-1/2 operators ${\vec{\tau}}_{j}$ which sit at the bond $(i,i+1)$ of the original lattice such that the states $|1_{i}1_{i+1}\rangle$ and $|0_{i}0_{i+1}\rangle$ correspond to $\tau_{j}^{z}=1$ and $\tau_{j}^{z}=-1$ respectively.
In the space of the degenerate ground states described above, we see that the $V_{2}$ term in Eq. (5) effectively acts as $(V_{2}/4)\tau_{j}^{z}\tau_{j+1}^{z}$. For instance, acting on the state $|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle$, the term $(V_{2}/4)(n_{i+1}^{f}-1/2)(n_{i+2}^{f}-1/2)$ gives $(V_{2}/4)|\tau_{j}^{z}=1,\tau_{j+1}^{z}=1\rangle$.
Next, the hopping term in Eq. (5) of the form $t(c_{i+1}^{\dagger}c_{i+2}+c_{i+2}^{\dagger}c_{i+1})$ flips the state $(n_{i},n_{i+1},n_{i+2},n_{i+3})$ from $(1,1,0,0)$ to $(1,0,1,0)$, and vice versa. (The first state has $\tau_{j}^{z}=1$ and $\tau_{j+1}^{z}=-1$. The second state does not have a description in terms of $\tau_{j}^{z}$, and it appears as an intermediate state in second-order perturbation theory). According to the unperturbed Hamiltonian, the second state has an energy which is higher than that of the first state by an amount equal to $V_{1}$. As a result, second-order perturbation theory gives us a term equal to $(1/2)(1-\tau_{j}^{z}\tau_{j+1}^{z})(t^{2}/V_{1})$.
Combining the terms of order $V_{2}$ and $t^{2}/V_{1}$, we find that the low-energy effective Hamiltonian is given by
$H_{eff}~=~\sum_{j}(\frac{V_{2}}{4}~-~\frac{t^{2}}{2V_{1}})~\tau_{j}^{z}\tau_{j+1}^{z},$ (60)
up to some constants. This describes the classical Ising model, and it has a first-order phase transition at $V_{2}=2t^{2}/V_{1}$. If $V_{2}<2t^{2}/V_{1}$, the ground states have either $\tau_{j}=+1$ for all $j$ (so that the particle configuration is $1111\cdots$) or $\tau_{j}=-1$ for all $j$ (then the particle configuration is $0000\cdots$). This is the PS phase. If $V_{2}>2t^{2}/V_{1}$, the ground states are $\tau_{j}=+1,-1,+1,-1,\cdots$ or $1,+1,-1,+1,\cdots$, so that the particle configuration is either $11001100\cdots$ or $00110011\cdots$. This is the CDW-II phase. The transition between the two phases is first-order because if $E_{0}(V_{1},V_{2})$ denotes the ground state energy and we hold $V_{1}$ fixed, then the derivative $dE_{0}/dV_{2}$ has a discontinuity when we cross the line $V_{2}=2t^{2}/V_{1}$.